OpenStudy (anonymous):

what is the solution set of the given equation 5x+2y=7 y=x+1

OpenStudy (anonymous):

well do you know how to do the substation property

OpenStudy (anonymous):

substitution

OpenStudy (anonymous):

no

OpenStudy (anonymous):

then you should go learn it because it is very important throughout algebra but I will show you 1 step and you do it ok?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

5x+2y=7 y=x+1 so you have to move the y= x + 1 into 2y 5x + 2(x + 1) = 7

OpenStudy (anonymous):

you should know it from there

OpenStudy (anonymous):

it will come out a demical right

OpenStudy (anonymous):

no

OpenStudy (anonymous):

I mean yes

OpenStudy (anonymous):

yay i got it thank you

OpenStudy (anonymous):

yw